suppose tom owns 2,000 shares of company X and 10,000 shares of company Y. The total value of tom's holdings of these two companies is $372,000.
Suppose tim owns 8,000 shares of company X and 6,000 of company Y. total value holdings is $400,000
write an equation for tom and tim's holdings. use variables X and Y to represent the value of shares of company X and company Y then need to show the value of share of company X and Y

Answer by Theo(13342)

Tom owns 2,000 shares of company X and 10,000 shares of company Y for a total value equal to $372,000.

Tim owns 8,000 shares of company X and 6,000 shares of company Y for a total value equal to $400,000.

Let a equal the value of each share of company X.

Let b equal the value of each share of company Y.

For Tom:

2,000 * a + 10,000 * b = 372,000.

For Tim:

8,000 * a + 6,000 * b = 400,000.

Since the value of each share in company X and company Y is the same, then a and b must be the same values in both equations.

To find that value, solve the 2 equations simultaneously.

The 2 equations that need to be solved simultaneously are:

2,000 * a + 10,000 * b = 372,000.
8,000 * a + 6,000 * b = 400,000.

Multiply the first equation by 4 to get:

8,000 * a + 40,000 * b = 1,488,000.
8,000 * a + 6,000 * b = 400,000

Subtract second equation from first equation to get:

34,000 * b = 1,088,000

Divide both sides of this equation by 44,000 to get:

b = 32 dollars per share.

Substitute for b in the first original equation to get:

2,000 * a + 10,000 * b = 372,000 becomes:

2,000 * a + 10,000 * 32 = 372,000

Solve for a to get:

a = 26 dollars per share.

You have:

a = 26 dollars per share.
b = 32 dollars per share.

Substitute in both original equations to get:

2,000 * a + 10,000 * b = 372,000.
8,000 * a + 6,000 * b = 400,000.

become:

2,000 * 26 + 10,000 * 32 = 52,000 + 320,000 = 372,000.
8,000 * 26 + 6,000 * 32 = 208,000 + 192,000 = 400,000.

Since the sum of the parts is equal to the original whole, this confirms the values are good.

Your answer is:

The value of the shares of company X is $26.00 per share.

The value fo the shares of company Y is $32.00 per share.